# MVPatch: More Vivid Patch for Adversarial Camouflaged Attacks on Object Detectors in the Physical World

### v2

@@ V-C1 Evaluation of Attack Performance of Diverse Adversarial Patches in the Physical Domain
Evaluation and Analysis for Experimental Consequence: As Table V illustrates, both AdvPatch(Fig.5) and AdvTetxure(Fig.5) do not have much more attack performance on the YOLOv2 model in the physical world and the ASRs are 13.57% and 12.72%, respectively. The Natural Patch(Fig.5) has higher ASR and its value of ASR is 19.43%, but the other Natural Patch(Fig.5) does not perform as well as Fig.5. The ASR of the first Natural Patch(Fig.5) is almost close to the Benign Image(Fig.5). The MVPatches perform better than other adversarial patches in the physical world. The ASR of MVPatches are respectively 22.60% (Fig.5) and 26.33% (Fig.5). Consequently, we ensure the MVPatch has more attack performance than others in the physical world.
V-C2 Evaluation of Attack Performance from Diverse Angles in the Physical Domain
@@
We collected 6846 images from various angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) to assess the impact of different angles on the attack performance of adversarial patches, including the Natural Patch, AdvPatch, AdvTexture, Benign Image, and our MVPatch. To keep the fairness of the experiment, we change angles from {$0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$} with the same distance (2 meters). The volunteer rotated while the adversarial patch rotated in the opposite direction, ensuring that the patch always faced the camera. We refer to this setup as ”with Rotation”. The experimental results are presented in Fig.6(a).
Explanation for Experimental Consequence: The light blue dotted line with squares represents the MVPatch with Rotation, while the red dotted line with circles represents the Benign Image with Rotation. The dark blue dotted line with triangles represents the AdvPatch with Rotation, while the green dotted line with triangles represents the AdvTexture with Rotation. The black dotted line with squares represents the Natural Patch with Rotation. The X-axis represents the angles, while the Y-axis represents the attack success rate.
@@
Evaluation and Analysis for Experimental Consequence: The effectiveness of adversarial patches with rotation diminishes as the angle of rotation increases. At an angle of $0^{\circ}$, the ASR of the MVPatch can reach 39.5%, and other patches perform well at the angle of $0^{\circ}$ except for the AdvTexture. At the angle of $90^{\circ}$, the ASRs of the Natural Patch and AdvTexture surpass the MVPatch at the same angle. Even a slight variation in the angle can significantly enhance the robustness and aggressiveness of the adversarial patch when it is rotated. Fig.6(a) also demonstrates that the MVPatch outperforms other attack methods in terms of angle variations (at the angles of $0^{\circ}$, $30^{\circ}$, and $60^{\circ}$).
V-C3 Evaluation of Attack Performance from Diverse Distance in the Physical Domain
A total of 3984 images are collected at various distances (1 meter, 2 meters, 3 meters, and 4 meters) to assess how different distances impact the attack performance of adversarial patches, specifically MVPatch, Natural Patch, AdvPatch, AdvTexture, and Benign Image. To keep the fairness of the experiment, we change distances from {1, 2, 3, 4} meters with the same angle ($0^{\circ}$). Fig.6(b) presents the results.
@@ V-C3 Evaluation of Attack Performance from Diverse Distance in the Physical Domain
Explanation for Experimental Consequence: MVPatch is represented by the light blue dotted line with squares. Natural Patch is represented by the black dotted line with squares. AdvPatch is represented by the dark blue dotted line with triangles. AdvTexture is represented by the green dotted line with triangles. The benign Image is represented by the red dotted line with circles. The X-axis represents the attack success rate, and the Y-axis represents the distance from the volunteer with the adversarial patch to the camera.
@@
Evaluation and Analysis for Experimental Consequence: As Fig.6(b) shows, the adversarial patches generated by MVPatch and AdvPatch exhibit high levels of aggression when the volunteer is at a distance of one meter from the camera. The ASR of MVPatch is 71.67% and the ASR of AdvPatch is 63.46% at a distance of 1 meter. As the volunteer moves farther away from the camera (beyond one meter), the effectiveness of the attack diminishes. However, the adversarial patches generated by AdvTexture and Natural Patch have higher aggression at a distance of 2 meters. When the volunteer moves farther away from the camera (beyond two meters), the effectiveness of the attack diminishes. Similarly, the attack weakens as the volunteer moves closer.
V-D Illustraton of Experimental Consequence
@@
This section provides experimental results in both the digital and physical domains to demonstrate our proposed approach. The green rectangle represents a successful detection of a person by the object detector, while the red rectangle indicates a failure to detect the person class. In the digital domain, we choose various experimental images from the Inria Person dataset as evaluation objects. In the physical domain, the source images of this experiment are not covered with mosaics. For privacy protection, mosaics are later added near the face. Fig.7 displays the results of digital experiments, while Fig.8 depicts the results of physical experiments.
VI Conclusion
@@
In this paper, we propose a dual-perception-based attack framework. This framework comprises a model-perception degradation method and a human-perception improvement method. The aim is to create vivid and aggressive adversarial patches, which we refer to as MVPatch, for person detection models suitable for real-world implementation. In terms of the model-perception degradation method, we exploit the ensemble attack on various object detectors to enhance the transferability of the adversarial patch. In terms of the human-perception improvement method, we introduce a lightweight approach for visual similarity measurement designed to make the adversarial patch less noticeable. Through extensive qualitative and quantitative experiments, we compare different approaches and find that minimizing the object loss yields the most effective patches while maintaining competitive attack performance compared to similar methods. We evaluate the transferability and naturalness of the crafted adversarial patches in both digital and physical domains. Extensive experiments demonstrate that the proposed MVPatch performs remarkable stealthiness and transferability with few computational costs.
However, there are still limitations in terms of factors such as angles and distance that can affect the performance of adversarial attacks. In future work, we aim to explore approaches to designing adversarial patches that can minimize the influence of angle and distance and maximize attack performance in both digital and physical domains.

### v3

@@ V-D1 Comparision with Other Adversarial Patches
Analysis and Discussion: As Table V illustrates, both AdvPatch(Fig.5) and AdvTetxure(Fig.5) do not have much more attack performance on the YOLOv2 model in the physical world and the ASRs are 13.57% and 12.72%, respectively. The Natural Patch(Fig.5) has higher ASR and its value of ASR is 19.43%, but the other Natural Patch(Fig.5) does not perform as well as Fig.5. The ASR of the first Natural Patch(Fig.5) is almost close to the Benign Image(Fig.5). The MVPatches perform better than other adversarial patches in the physical world. The ASR of MVPatches are respectively 22.60% (Fig.5) and 26.33% (Fig.5). Consequently, we ensure the MVPatch has more attack performance than others in the physical world.
V-D2 Diverse Angles
@@
We collected 6,846 images from various angles ($0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$) to assess the impact of angle variation on the attack performance of different adversarial patches, including Natural Patch, AdvPatch, AdvTexture, Benign Image, and our MVPatch. To ensure experimental fairness, angles were adjusted to $0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $90^{\circ}$ while maintaining a constant distance of 2 meters. The volunteer rotated, with the adversarial patch rotating in the opposite direction to ensure it always faced the camera, a setup referred to as ”with Rotation.” The experimental results are presented in Fig.6(a). The light blue dotted line with squares represents the MVPatch with Rotation; the red dotted line with circles represents the Benign Image with Rotation; the dark blue dotted line with triangles represents the AdvPatch with Rotation; the green dotted line with triangles represents the AdvTexture with Rotation; and the black dotted line with squares represents the Natural Patch with Rotation. The X-axis represents the angles, while the Y-axis represents the attack success rate.
@@
Analysis and Discussion: The effectiveness of adversarial patches with rotation diminishes as the angle of rotation increases. At an angle of $0^{\circ}$, the ASR of the MVPatch can reach 39.5%, and other patches perform well at the angle of $0^{\circ}$ except for the AdvTexture. At the angle of $90^{\circ}$, the ASRs of the Natural Patch and AdvTexture surpass the MVPatch at the same angle. Even a slight variation in the angle can significantly enhance the robustness and aggressiveness of the adversarial patch when it is rotated. Fig.6(a) also demonstrates that the MVPatch outperforms other attack methods in terms of angle variations (at the angles of $0^{\circ}$, $30^{\circ}$, and $60^{\circ}$).
V-D3 Diverse Distance
@@
A total of 3,984 images were collected at various distances (1 meter, 2 meters, 3 meters, and 4 meters) to evaluate the impact of distance on the attack performance of adversarial patches, including MVPatch, Natural Patch, AdvPatch, AdvTexture, and Benign Image. To ensure experimental fairness, distances were varied 1, 2, 3, 4 meters while maintaining a constant angle of $0^{\circ}$. The results are presented in Fig.6(b). The light blue dotted line with squares represents the MVPatch, the black dotted line with squares represents the Natural Patch, the dark blue dotted line with triangles represents the AdvPatch, the green dotted line with triangles represents the AdvTexture, and the red dotted line with circles represents the Benign Image. The X-axis denotes the attack success rate, and the Y-axis indicates the distance from the volunteer with the adversarial patch to the camera.
@@
Analysis and Discussion: As Fig.6(b) shows, the adversarial patches generated by MVPatch and AdvPatch exhibit high levels of aggression when the volunteer is at a distance of one meter from the camera. The ASR of MVPatch is 71.67% and the ASR of AdvPatch is 63.46% at a distance of 1 meter. As the volunteer moves farther away from the camera (beyond one meter), the effectiveness of the attack diminishes. However, the adversarial patches generated by AdvTexture and Natural Patch have higher aggression at a distance of 2 meters. When the volunteer moves farther away from the camera (beyond two meters), the effectiveness of the attack diminishes. Similarly, the attack weakens as the volunteer moves closer.
V-E Visualization
@@
This section provides experimental results in both the digital and physical domains to visualize our proposed approach. The green rectangle represents a successful detection of a person by the object detector, while the red rectangle indicates a failure to detect the person class. In the digital domain, we choose various experimental images from the Inria Person dataset as evaluation objects. In the physical domain, the source images of this experiment are not covered with mosaics. For privacy protection, mosaics are later added near the face. Fig.7 displays the results of digital experiments, while Fig.8 depicts the results of physical experiments.
VI Conclusion and Limitations
@@
In this paper, we propose a Dual-Perception-Based Framework (DPBF) that includes a Model-Perception-Based Module (MPBM), a Human-Perception-Based Module (HPBM), and regularization terms. Our objective is to create vivid and aggressive adversarial patches, termed More Vivid Patch (MVPatch), designed to camouflage individuals in detection models for real-world applicability. The MPBM leverages ensemble attacks across various object detectors to enhance the transferability of adversarial patches, demonstrating improved generalization and stability through generalization error analysis. Meanwhile, the HPBM introduces a lightweight approach for visual similarity measurement, which makes the adversarial patches less noticeable and incorporates additional transformations to further enhance transferability while preserving stealthiness. Regularization terms are employed to increase the practicality of the generated patches in physical environments. We also introduce metrics for naturalness and transferability to provide an unbiased assessment of the adversarial patches. Extensive qualitative and quantitative experiments reveal that minimizing our proposed objective function results in the most effective patches, maintaining competitive performance compared to existing methods. The MVPatch demonstrates exceptional stealthiness and transferability across both digital and physical domains, achieving these results with minimal computational cost.
Limitations. Despite the advancements presented, our approach has several limitations. Factors such as angles and distances can still affect the effectiveness of adversarial attacks. Additionally, the vulnerability of adversarial patches against state-of-the-art object detection models remains inadequately explored. Future work will address these issues by developing methods to minimize the impact of angle and distance, thereby optimizing attack performance in both digital and physical contexts. We also intend to apply our technique to a wider range of off-the-shelf object detection models to evaluate their susceptibility to adversarial patches. Furthermore, we plan to explore more robust defense mechanisms to mitigate the risk of adversarial attacks in both digital and physical environments.